A company would like to classify its costs according to cost behavior. The following data is available about one of its costs:

Q1

Q2

Q3

Production in Units

6,000

7,000

5,800

Cost

120,000

140,000

116,000

The cost behavior is best described as

  

a. None of above

   

b. Variable Cost

   

c. Fixed Cost

   

d. Mixed Cost
Answer #1

Cost per unit at different levels of production

6,000 units = 120,000/6,000 = 20

7,000 units = 140,000/7,000 = 20

5,800 units = 116,000/5,800 = 20

As the cost per unit is fixed at different levels of product, the cost is variable in nature.

Option B is the answer